You know you will have $500,000 in your account when you retire. You want to be able to take monthly withdrawals from the account for a total of 30 years. Your retirement account earns 8% interest. How much will you be able to withdraw each month?


1
Expert's answer
2021-10-12T09:03:24-0400

"\displaystyle
P = \frac{w(1-(1+\frac{r}{k})^{-kt}}{\frac{r}{k}}\\
\text{where P = balance at the beginning of the period}\\
\text{w = regular withdrawal rate}\\
\text{r = annual interest rate}\\
\text{k = no of compounding periods in one year}\\
\therefore 500000 = \frac{w(1-(1+\frac{0.08}{12})^{-360}}{\frac{0.08}{12}}\\
\implies w = \$3667.03"